Emile receives a flat weekly salary of $240 plus %12 commission of the total volume of all sales he makes.What must his dollar volume be in a week if he is to make a total weekly salary of $540?
Answer by checkley77(12844)

240+.12x=540
.12x=540-240
.12x=300
x=300/.12
x=2,500 is the required in sales for the week.
Proof:
240+.12*2,500=540
240+300=540
540=540